OpenStudy (aaronandyson):

The area of a circle is 64 pi sq.cm.Find the length of its arch subtending all the angles of 60 degrees. Also find the area of the corresponding sector

OpenStudy (aaronandyson):

@rebeccaxhawaii

OpenStudy (aaronandyson):

@Qwertty123

OpenStudy (qwertty123):

so..a = 1/2 r^2 θ, 1/2 (60^2) θ = 64 pi? Not sure if I did that right..@.Sam.

OpenStudy (qwertty123):

@.Sam.

OpenStudy (aaronandyson):

60 is not the radius

OpenStudy (aaronandyson):

60 is the angle

OpenStudy (aaronandyson):

\[\frac{ 60 }{ 360 }\pi r^2\]

OpenStudy (aaronandyson):

thats 100.48/3

OpenStudy (qwertty123):

whoops

OpenStudy (qwertty123):

what is 100.48 divided by 3?

OpenStudy (aaronandyson):

whats the length of the arch?